esolving gel with DI water, and blot dry with a paper towel. (demo) a small flask (clean and reuse the old flask), mix com , Not sure how to figure out what formulas to use here
Answer #1

Total volume of the gel = 5 uL
Water = 3.4 mL

0.83 mL of 30% acrylamide mix:
1 mL in 5 mL = 1/5 = 0.2
0.83 mL in 5 mL = 0.83/5 = 0.166
Final Concentration = 0.166*30 = 4.98% =~5%

0.63 mL of 1 M Tris:
0.83 mL in 5 mL = 0.63/5 = 0.126
Final Concentration = 0.126*1 = 0.126 M

0.05 mL of 10% SDS:
0.05 mL in 5 mL = 0.05/5 = 0.01
Final Concentration = 0.01*10 = 0.1%

0.05 mL of 10% APS:
0.05 mL in 5 mL = 0.05/5 = 0.01
Final Concentration = 0.01*10 = 0.1%

5 uL = 0.005 mL of 100% TEMED:
0.005 mL in 5 mL = 0.005/5 = 0.001
Final Concentration = 0.001*10 = 0.01%